OpenStudy (lgbasallote):

@amistre64 can you tell me how to use your method in exact DE for this?? \[(2x+y)dx + (x^2 - x)dy = 0\] it seems i end up in a dead end :/

OpenStudy (anonymous):

linear DE! i got integration factor=(x-1)/x

OpenStudy (experimentx):

\[ e^{\int{\frac1{x(x-1)}}dx}\]

OpenStudy (anonymous):

the ans is \[y(x-1/x)+logx=c\] @experimentX is it correct?

OpenStudy (anonymous):

(x-1)/x not x-1/x

OpenStudy (experimentx):

lol not sure ... \[ \frac{1}{x(x-1)} = \frac{-1}{x} + \frac{1}{x-1}\]

OpenStudy (experimentx):

Looks like you are right!!

OpenStudy (anonymous):

hmn!

OpenStudy (amistre64):

does Fxy = Fyx?

OpenStudy (amistre64):

2x NOTequal 2x-1; therefore its not "exact"

OpenStudy (amistre64):

opps, well, same unequal outcome but i didnt do the left part right did i?

OpenStudy (amistre64):

2x + y wrt y = 1 x^2-x wrt x = 2x -1 1 notequal 2x-1 so it aint exact

OpenStudy (lgbasallote):

so it's not exact?? weird :/